# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v1

@@ VI No-go theorems for 1D decoding graphs of dynamical codes
Let us consider the implications of Theorem 2. In static codes, as we have pointed out, this repetition code structure in the decoding problem can lead to $50\%$ code capacity thresholds under infinitely biased noise. In GDCs, we do not expect to see such high thresholds owing to Theorem 2. The reason that static codes perform better than dynamical codes under biased noise is that they obey two separate symmetries under an infinitely-biased code-capacity noise model. There is a spatial symmetry that forces syndromes to appear in pairs along 1D strips of the lattice, but there is also a temporal symmetry: syndromes must appear in pairs within a given time slice. This is because a code-capacity noise model includes no measurement errors. These two symmetries simplify the detector graph into a collection of disjoint repetition code decoding graphs, leading to a high performance of these codes under biased noise. GDCs necessarily break the temporal symmetry (for HDCs, this is a consequence of Lemma 1), leading to edges between vertices in different time slices of the detector graph even without measurement errors (see Fig. 3). We thus expect such codes can only possess a single symmetry under an infinitely-biased code-capacity noise model, as the X3Z3 Floquet code does.
@@ VII Conclusions and outlook
In this paper, we introduce the X3Z3 Floquet code, the first bias-tailored dynamical code based on two-qubit parity check measurements. We show that, despite having no constant stabilisers, the code has a persistent symmetry under pure dephasing (or pure bit-flip) noise which allows for simplified decoding. This results in a substantially improved threshold and sub-threshold performance under biased noise, when compared to other Floquet codes. We demonstrate the enhanced performance through our simulation results obtained from using a fast matching decoder and a noise model approaching realistic circuit-level noise. Besides the superior performance of the X3Z3 Floquet code, our results also demonstrate that there are differences, albeit modest, in the threshold and subthreshold performance of the XYZ2 and P6 honeycomb codes under biased noise.
To explain why the X3Z3 Floquet code does not reach the same high performance as the bias-tailored static codes, we prove that a dynamical code on the honeycomb lattice (obeying certain assumptions common to standard Floquet codes) cannot have a so-called 1D-decodable $Z$-detector graph, the crucial requirement for the high performance of static codes (at least those without hyperedge-like syndromes). Despite this limitation, the bias-tailored X3Z3 Floquet code has the advantage over its static counterpart in that it requires only lower-weight measurements. Specifically for devices without native two-qubit parity measurements, we devise bias-preserving parity check measurement circuits that allow for high implementation of the code and can be applied to any qubit type. Our work therefore demonstrates that the X3Z3 Floquet code is a leading quantum error correction code especially for devices with limited connectivity such as the hexagonal and heavy-hexagonal architectures.
We now give several directions for future work. While in this paper we use multi-Pauli product gates for the parity check measurements which are native to certain architectures (for example, Majorana qubits) we leave for future work the study of the code performance using the bias-preserving parity measurement circuits proposed in Sec. III.2. Furthermore, it would also be interesting to investigate the performance of the X3Z3 Floquet code using a better decoder (such as a tensor network decoder) and to consider ways of analytically deriving the best achievable thresholds for the X3Z3 Floquet code.
@@ Acknowledgements.
We would like to thank our colleagues at Riverlane, especially Alexandra Moylett, Ophelia Crawford, and György Gehér for helpful discussions. We are also grateful to Ophelia Crawford for her reading and comments on the manuscript. CM would like to thank Benjamin Brown for pointing out a useful circuit for performing next-nearest-neighbour entangling gates.
@@ A.1 Honeycomb code
We here provide details of the two honeycomb code variants discussed in the main text. The codes’ logical operators evolve through their measurement cycles. In Fig. 9(a), we show one set of logical operators evolving through two measurement cycles. While the measurement sequence has period 3, the logical operators only return to their initial values (up to signs) with period 6.
In Fig. 9(b) we show the plaquette and edge operators for both honeycomb code variants. The P6 code has edge operators with Pauli bases defined by the edge colour, while the XYZ2 code has edge operators with bases defined by the edge’s orientation and position relative to neighbouring plaquettes [see Fig. 9(b)]. The plaquette operators of the P6 code also differ between plaquettes of different colours, while the plaquette operators of the XYZ2 code are all the same. These plaquette operators are products of edge operators along the boundary of the plaquette. For example, a red plaquette operator in the P6 code is the product of the green and blue $YY$ and $ZZ$ (respectively) edge operators around its boundary. We measure plaquette operators in two rounds. For example, after measuring $g\rightarrow b$ checks, we can learn the value of an $X$ plaquette by multiplying the measurement values of green and blue checks around its boundary. Unlike the CSS and X3Z3 Floquet codes, the honeycomb code has constant plaquette stabilisers.
@@ A.2 CSS Floquet code
In Fig. 10, we illustrate the measurement cycle of the CSS Floquet code and the resulting evolution of (one set of) its logical operators. Its logical operators are either of $X$-type or $Z$-type for all time, as shown. We show an $X$-type logical operator defined on a vertical non-trivial cycle of the lattice. However, there exists another $Z$-type logical operator (not shown) defined on that same vertical cycle. One can find this operator (and its anti-commuting partner) by noting the symmetry in the code obtained by proceeding forward by three measurement subrounds ($t\mapsto t+3$) and applying a Hadamard gate to all qubits ($X\leftrightarrow Z$).
The plaquette operator members of the ISG at each step of the measurement cycle are indicated in Fig. 10. As can be seen, after each measurement subround, plaquettes of one colour host only a single ($X$- or $Z$-type) member of the ISG. To consider detectors, let us follow a single blue $X$ plaquette through the measurement sequence. After red $X$ checks (top left of Fig. 10), the plaquette operator is measured (we infer its outcome by multiplying together the check measurement outcomes around its boundary). This initialises the plaquette. It is only then re-measured after green $X$ check measurements, at which point a detector is formed. After this, blue $Z$ checks are measured, which anti-commute with the plaquette operator and its subsequent measurement outcomes are random.
Figure 9: The honeycomb code measurement cycle, logical operators, plaquettes and edges. (a) A distance-4 (under code-capacity noise) P6 code [7] is shown throughout its evolution. Edge check measurements are defined in (b). One set of anti-commuting logical operators is shown: a vertical logical defined on the yellow-shaded non-trivial cycle of the torus, and a horizontal logical defined on the grey-shaded non-trivial cycle. The logicals are defined as the product of Pauli operators on coloured qubits in the cycle, where red, green and blue qubits support $X$, $Y$, and $Z$ Pauli bases, respectively. The other set of logical operators is offset by three measurement subrounds. (b) Top: The plaquette operators of the P6 honeycomb code are shown along with the edge operators measured (red, green and blue edges/plaquettes support $X$, $Y$ and $Z$ operators, respectively). Bottom: All plaquette operators of the XYZ2 honeycomb model are the same and are shown on the left. Edge operators are defined based on their orientation in the T junction (rather than edge colour), as shown on the right.
Figure 10: Measurement cycle for the CSS Floquet code. $X$ or $Z$ edge operators are measured in each time step, with the colour of edge and Pauli type indicated by the labelled arrows. Plaquettes labelled with $X$ or $Z$ labels host only a stabiliser operator of the corresponding type, while other plaquettes host both $X$- and $Z$-type stabilisers. One set of anti-commuting logical operators is shown.
@@ Appendix B Noise models
In this paper, we consider two different noise models: code-capacity and entangling measurement (SDEM3) noise models. In the code-capacity noise model, we apply single-qubit Pauli noise on all the data qubits independently at every measurement subround. The single-qubit Pauli noise channel is given by
@@
Besides the code-capacity noise model, in this paper we also use the SDEM3 noise model which assume noisy multi-qubit Pauli product measurement (MPP) gates. The SDEM3 noise model applies single-qubit noise after every single-qubit gate, reset and measurement, two-qubit noise after each MPP gate, and a random classical flip after every measurement. We assume each of the error channels occurs with a total probability $p$, as in Ref. LABEL:Gidney_Planar_honeycomb_2022. To take into account biased noise, we generalise the SDEM3 depolarising noise model used in Refs. [9, 18]. Here the single-qubit noise channel has the same form as given in Eq. (2) for the code-capacity noise model.
On top of the single-qubit errors, this model also has a two-qubit noise channel given by
$\mathcal{E}_{2q}(\rho)=(1-p)\rho+\sum_{O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}}p_{O}\hat{O}\rho\hat{O},$ (3)
which is applied after each of the MPP gates. To conform with how the bias is defined for the single-qubit noise channels, we also use $\eta$ to characterize the bias of the two-qubit noise channel. Specifically, we define the bias $\eta$ for the two-qubit noise channel such that
@@
Just as for the single-qubit noise, the above definition of $\eta$ ensures that the two-qubit noise at three special points $\eta=0$, $\eta=0.5$ and $\eta=\infty$ has the required forms (no $Z$ errors, depolarising, and pure dephasing, respectively). Given that the two-qubit error probabilities at these $\eta$ values must satisfy the conditions above, we define $\zeta\in[0,1]$ and write the two-qubit Pauli error probabilities as
@@
Note that Eq. (B) is defined such that the total probability of all the Pauli errors $\sum_{O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}}p_{O}=p$ regardless of the noise bias $\eta$. Apart from the single- and two-qubit Pauli noise channels, we also apply a classical flip of the measurement results with probability $p$ after each of the single and two-qubit measurements. These measurement flips are uncorrelated with the single and two-qubit noise channels.
@@ Appendix C Details of numerical simulations
For each distance $d_{\mathrm{eff}}=d/2$ (where $d_{\mathrm{eff}}$ and $d$ are the minimum number of fault locations required to produce a logical error for the SDEM3 and code-capacity depolarising noise models, respectively), we construct a $2d_{\mathrm{eff}}\times 3d_{\mathrm{eff}}$ honeycomb lattice with periodic boundary conditions. We run quantum memory experiments for $3d_{\mathrm{eff}}$ QEC rounds or $18d_{\mathrm{eff}}$ measurement subrounds (1 QEC round consists of 6 measurement subrounds). We choose a single logical qubit for each Floquet code (they all encode two logical qubits on a lattice with periodic boundary conditions) to test for the logical failure probability. Depending on the codes and parameter regimes, we run different numbers of Monte Carlo shots ranging from $10^{6}-10^{10}$ for each of the horizontal and vertical logical observables, which give us the horizontal $p_{H}$ and vertical $p_{V}$ logical failure probabilities. We then report the combined logical error probability:
@@
which is an estimate of the probability that either a vertical or a horizontal logical error occurs. Equation (6) assumes that the horizontal and vertical logical errors occur independently. We run the simulations of the four different Floquet codes for different values of physical error rates $p$ with various strengths of noise bias $\eta$ and under the two different noise models.

### v2

@@ VI No-go theorems for 1D decoding graphs of dynamical codes
The proof of the theorem is given in Appendix F. Owing to this theorem, we do not expect to see thresholds in dynamical codes as high as 50%. The reason that static codes perform better than dynamical codes under biased noise is that they obey two separate symmetries under an infinitely-biased code-capacity noise model. There is a spatial symmetry that forces syndromes to appear in pairs along 1D strips of the lattice, but there is also a temporal symmetry: syndromes must appear in pairs within a given time slice. These two symmetries simplify the detector graph into a collection of disjoint repetition code decoding graphs, leading to a high performance (up to 50% threshold) of the static codes under biased noise. On the other hand, even for the case where measurement errors are not considered, GDCs necessarily break the temporal symmetry (for HDCs, this is a consequence of Lemma 1 in Appendix F), leading to edges between vertices in different time slices of the detector graph (see Fig. 3). We thus expect that such codes can only possess a single symmetry under an infinitely-biased code-capacity noise model, as the X3Z3 Floquet code does.
@@ VII Conclusions and outlook
In this paper, we introduce the X3Z3 Floquet code, the first bias-tailored dynamical code based on two-qubit parity check measurements. We show that, despite having no constant stabilisers, the code possesses a persistent symmetry under pure dephasing (or pure bit-flip) noise which allows for a simplified decoding. This results in a substantially improved threshold and sub-threshold performance under biased noise, when compared to other Floquet codes. We demonstrate the enhanced performance through our simulation results obtained from using a fast matching decoder and two error models: a simplistic code-capacity noise model and a noise model approaching realistic circuit-level errors. Besides the superior performance of the X3Z3 Floquet code, our results also show that there are differences, albeit modest, in the threshold and subthreshold performance between the XYZ2 and P6 honeycomb codes under biased noise.
To explain why the X3Z3 Floquet code does not reach the same high performance as the bias-tailored static codes, we prove that a dynamical code on the honeycomb lattice (obeying certain assumptions common to standard Floquet codes) cannot have a so-called 1D-decodable $Z$-detector graph, the crucial requirement for the high performance of static codes. Despite this limitation, the bias-tailored X3Z3 Floquet code has the advantage over its static counterpart in that it requires only lower-weight measurements. Specifically for devices without native two-qubit parity measurements, we devise bias-preserving parity check measurement circuits for any qubit architecture, which allow for high-performance implementation of the code. Our work therefore demonstrates that the X3Z3 Floquet code is a leading quantum error correction code especially for devices with limited connectivity such as the hexagonal and heavy-hexagonal architectures.
We now give several directions for future work. While in this paper we use multi-Pauli product gates for the parity check measurements which are native to certain architectures (for example, Majorana [56, 10, 57] and photonic [58, 17] platforms) we leave for future work the study of the code performance using the bias-preserving parity measurement circuits proposed in Sec. III.2. Furthermore, it would also be interesting to investigate the performance of the X3Z3 Floquet code using a better decoder (such as a tensor network decoder) and to consider ways to analytically derive the best achievable thresholds for the X3Z3 Floquet code.
Author Contributions
FS developed the X3Z3 Floquet code and biased noise models. FS also performed the numerical simulations for all Floquet codes investigated. CM characterised the symmetry and decoding features of the codes tested. In addition, CM developed the bias-tailored circuits and the proofs of the no-go theorems. Both authors contributed to the writing of the manuscript.
@@ Acknowledgements.
We would like to thank our colleagues at Riverlane, especially Alexandra Moylett, Ophelia Crawford, and György Gehér for helpful discussions. We are also grateful to Ophelia Crawford and Hari Krovi for their comments on the manuscript. CM would like to thank Benjamin Brown for pointing out a useful circuit for performing next-nearest-neighbour entangling gates.
@@ A.1 Honeycomb code
We here provide details of the two honeycomb code variants discussed in the main text. The codes’ logical operators evolve through their measurement cycles. In Fig. 10(a), we show one set of logical operators evolving through two measurement cycles. While the measurement sequence has a period of 3, the logical operators only return to their initial values (up to signs) with a period of 6.
In Fig. 10(b), we show the plaquette and edge operators for both honeycomb code variants. The P6 code [8] has edge operators with Pauli bases defined by the edge colour, while the XYZ2 code [6] has edge operators with bases defined by the edge’s orientation in the T junction, i.e., horizontal for the $Z$ edges, and rotated 90 clockwise and counter-clockwise from the horizontal edges, for the $X$ and $Y$ edges, respectively [see Fig. 10(b)]. The plaquette operators of the P6 code also differ between plaquettes of different colours, while the plaquette operators of the XYZ2 code are all the same. Each plaquette operator is a product of the edge operators along the plaquette boundary. For example, a red plaquette operator is the product of the green and blue edge operators around its boundary. The plaquette operators are measured in two subrounds. For instance, after measuring the $g$ and $b$ checks sequentially, we can learn the value of each red plaquette by multiplying the measurement values of the green and blue checks around its boundary. Unlike the CSS and X3Z3 Floquet codes, the honeycomb code has constant plaquette stabilisers.
Figure 10: The honeycomb code measurement cycle, logical operators, plaquettes and edges. (a) A distance-4 (under code-capacity noise) P6 code [8] is shown throughout its evolution. The type of check measured at each measurement subround is indicated by the arrows and the highlighted/coloured edges in the lattice. One set of anti-commuting logical operators is shown in yellow and brown-coloured paths. Each logical is a product of Pauli operators on the coloured qubits along its cycle, where red, green and blue qubits support $X$, $Y$, and $Z$ Pauli bases, respectively. The other set of logical operators is offset by three measurement subrounds. (b) Top: Red, green and blue edges/plaquettes of the P6 honeycomb code support $X$, $Y$ and $Z$ operators, respectively. Bottom: All plaquettes of the XYZ2 honeycomb code support the same XYZ2 operator, as shown on the left. Edge operators are defined based on their orientation in the T junction (rather than the edge colour), as shown on the right.
@@ A.2 CSS Floquet code
In Fig. 11, we illustrate the measurement cycle of the CSS Floquet code [18, 13] and the resulting evolution of (one set of) its logical operators. Its logical operators are either of $Z$-type or $X$-type for all time, as shown. We show a $Z$-type logical operator defined on a vertical non-trivial cycle of the lattice. However, there exists another $X$-type logical operator (not shown) defined on that same vertical cycle. One can find this operator (and its anti-commuting partner) by noting the symmetry in the code obtained by proceeding forward by three measurement subrounds ($t\mapsto t+3$) and applying a Hadamard gate to all qubits ($X\leftrightarrow Z$).
The plaquette operator members of the ISG at each step of the measurement cycle are indicated in Fig. 11. The value of each plaquette operator is inferred, in one time step, from the product of the measurement values of three edges surrounding the plaquette. At each measurement subround, only one kind [which is of one type ($X$ or $Z$) and one colour] of plaquette is reinitialised and one other kind is re-measured to form a detector. The reinitialised and re-measured plaquettes have the same ($X$ or $Z$) type as the checks performed at that particular subround, but are of different colours from each other and also from the checks [18, 13]. The colours of the reinitiliased and re-measured plaquettes are respectively the same as the colours of the checks performed just before and immediately after the current subround.
To see how detectors are formed in this code, let us follow a single blue $X$ plaquette through the measurement sequence. After red $X$ checks (top left of Fig. 11), the blue $X$ plaquette operator is measured (we infer its outcome by multiplying together the check measurement outcomes around its boundary). This initialises the plaquette. It is only re-measured after the green $X$ check measurements, four subrounds later, at which point a detector is formed. After this, blue $Z$ checks are measured, which anti-commute with the blue $X$ plaquette operator. Hence, its measurement outcome after this subround is random.
Figure 11: Measurement cycle for the CSS Floquet code. $X$ or $Z$ edge operators are measured alternately, with the colour of edge and Pauli type indicated by the labelled arrows. Plaquettes labelled with $X$ or $Z$ labels host only a stabiliser operator of the corresponding type, while other plaquettes host both $X$- and $Z$-type stabilisers. One set of anti-commuting logical operators is shown by yellow and brown paths. Each logical is a product of only one type of Pauli operators, either $X$ or $Z$, represented by the red or blue coloured qubits.
@@ Appendix B Noise models
Figure 12: Logical failure probability $p_{L}$ of the X3Z3 Floquet code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}-10^{7}$ number of shots.
In this paper, we consider two noise models: code-capacity and entangling measurement (SDEM3) noise models. In the code-capacity noise model, we apply single-qubit Pauli noise on all the data qubits independently at every measurement subround. The single-qubit Pauli noise channel is given by
@@
Besides the code-capacity noise model, in this paper we also use the SDEM3 noise model which assumes noisy multi-Pauli product (MPP) measurement gates. The SDEM3 noise model applies single-qubit noise after every single-qubit gate, reset and measurement, two-qubit noise after each MPP gate, and a random classical flip after every measurement. We assume each of the error channels occurs with a total probability $p$, as in Ref. [9]. To take into account biased noise, we generalise the SDEM3 model for depolarising noise used in Refs. [9, 15]. Here the single-qubit noise channel has the same form as given in Eq. (2) for the code-capacity noise model.
Figure 13: Logical failure probability $p_{L}$ of the CSS Floquet code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}$ number of shots.
On top of the single-qubit errors, this model also has a two-qubit noise channel applied after each of the MPP gates, which is given by
$\mathcal{E}_{2q}(\rho)=(1-p)\rho+\sum_{O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}}p_{O}\hat{O}\rho\hat{O}.$ (3)
To conform with how the bias is defined for the single-qubit noise channel, we also use $\eta$ to characterize the bias of the two-qubit noise channel. Specifically, we define the bias $\eta$ for the two-qubit noise channel such that
@@
Just as for the single-qubit noise, the above definition of $\eta$ ensures that the two-qubit noise at the three special points, $\eta=0$, $\eta=0.5$ and $\eta=\infty$, are $Z$-error free, depolarising, and pure dephasing, respectively. Given that the two-qubit error probabilities at these $\eta$ values must satisfy the conditions above, we define $\zeta\in[0,1]$ and write the two-qubit Pauli error probabilities as
@@
Note that Eq. (B) is defined such that the total probability of all the Pauli errors $\sum_{O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}}p_{O}=p$ for all noise biases $\eta$. Apart from the single- and two-qubit Pauli noise channels, we also apply a classical flip of the measurement results with probability $p$ after each of the single and two-qubit measurements. These measurement flips are uncorrelated with the single and two-qubit noise channels.
@@ Appendix C Details of numerical simulations
Figure 14: Logical failure probability $p_{L}$ of the P6 honeycomb code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}$ number of shots.
For each distance $d_{\mathrm{eff}}=d/2$ (where $d_{\mathrm{eff}}$ and $d$ are the minimum number of fault locations required to produce a logical error for the SDEM3 and code-capacity depolarising noise models, respectively), we construct a $2d_{\mathrm{eff}}\times 3d_{\mathrm{eff}}$ honeycomb lattice with periodic boundary conditions. We run quantum memory experiments for $3d_{\mathrm{eff}}$ QEC rounds or $18d_{\mathrm{eff}}$ measurement subrounds (1 QEC round consists of 6 measurement subrounds). The simulations are performed for four different Floquet codes using different values of physical error rates $p$ and various strengths of noise bias $\eta$. Moreover, we consider two different noise models: code-capacity and SDEM3 noise models as detailed in Appendix B.
We choose a single logical qubit for each Floquet code (they all encode two logical qubits on a lattice with periodic boundary conditions) to test for the logical failure probability. Depending on the codes and parameter regimes, we run different numbers of Monte Carlo shots ranging from $10^{5}-10^{10}$ (larger number of shots for longer code distances and smaller physical error rates) for each of the horizontal and vertical logical observables, which give us the horizontal $p_{H}$ and vertical $p_{V}$ logical failure probabilities. We then report the combined logical error probability:
@@
which is an estimate of the probability that either a vertical or a horizontal logical error occurs. Equation (6) assumes that the horizontal and vertical logical errors occur independently. For the infinitely biased code-capacity noise model, one of the logical errors which is of pure $Z$-type or pure $X$-type has zero logical error probability, because the pure biased noise cannot form the other logical operator which is of mixed-type and hence cannot flip the pure-type logical operator. As a result, for the infinitely biased code-capacity noise, the maximum combined logical error probability is $p_{L}=0.5$ [Eq. (6)]. This is in contrast to other noise biases and models where the maximum combined logical error probability is $p_{L}=0.75$ since $p_{V},p_{H}\leq 0.5$.